OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER! Estimate the circumference of the circle with the given radius or diameter. use 3.14 for n. Round to the nearest unit. A-27 in. B- 79 in. C- 1,809 in. D- 151 in.

OpenStudy (anonymous):



OpenStudy (anonymous):

@Mathbreaker can you help?

OpenStudy (anonymous):

Circumference is always equal to 2*pi*r hence when r = 24 2*3.14*24 = circumference = 3.14*48 = 150.72 which is 151 inches

OpenStudy (anonymous):

thanks you so much!

OpenStudy (anonymous):

Youre welcome :)